can you please explain what is kirchoff’s law in currentelectricity

Grade:10

1 Answers

hero
6 years ago

This law is also called Kirchhoff's first lawKirchhoff's point rule, or Kirchhoff's junction rule (or nodal rule).

The principle of conservation of electric charge implies that:

 flowing into that node is equal to the sum of currents flowing out of that nodecurrents, the sum of electrical circuitAt any node (junction) in an 

or equivalently

{\displaystyle \sum _{k=1}^{n}V_{k}=0}\sum _{k=1}^{n}V_{k}=0

Similar to KCL, it can be stated as:

 available in that loop.emfThe algebraic sum of the products of the resistances of the conductors and the currents in them in a closed loop is equal to the total   in any closed loop is equivalent to the sum of the potential drops in that loop, or: emfsMore simply, the sum of the   (voltage) around any closed network is zero, or: potential differencesThe directed sum of the electrical 

The principle of conservation of energy implies that

This law is also called Kirchhoff's second lawKirchhoff's loop (or mesh) rule, and Kirchhoff's second rule.

The algebraic sum of currents in a network of conductors meeting at a point is zero.